# Nudged-Elastic Band Calculations of Polymorph Transitions and Solid-State Reactions in Molecular Crystals

### v1

@@ Abstract
The modeling of solid-state transformations, such as polymorphic transitions and chemical reactions in molecular crystals, is vital for many applications like drug design or the development of new synthesis methods but faces several crucial challenges. First, the automatic initial pathway generation in terms of nudged-elastic band (NEB) calculations based on a linear interpolation often fails for periodic systems, leading to unrealistic geometries and atomic collisions. Second, the necessary system sizes are typically beyond the scope of density functional theory (DFT) calculations in terms of computational cost, but the associated accuracy is vitally needed. To address these issues, we introduce a novel hybrid interpolation method that combines linear interpolation for cell parameters with spherical linear interpolation (SLERP) for molecular structures or intramolecular fragments, ensuring smooth and realistic transitions. In addition, we also benchmark the accuracy of the recently introduced SO3krates machine-learned force field (MLFF) architecture in this context, which has the potential to substantially improve the computational efficiency. We apply these methods to two polymorph transitions and a solid-state Diels-Alder reaction and show the applicability of our interpolation method. Furthermore, we find excellent agreement of MLFF results with DFT energy barriers and pathways, when trained on appropriate reference data, potentially paving the way for more automated explorations of solid-state transformations in molecular crystals.
@@ I Introduction
Understanding solid-state transformations like transitions between molecular crystal polymorphs or solid-state chemical reactions at a molecular level is vital for pharmaceuticals, chemistry, and materials scienceKaupp (2002); Datta and Grant (2004); Anthony (2006); Saha, Nath, and Thakuria (2022). Polymorphs are different crystal-packing arrangements of the same molecule and although the chemical composition of a molecular crystal does not change when one polymorph transforms into another, this can often be associated with quite drastic changes in several properties. One important example crucial for the pharmaceutical industry is that polymorph transitions can affect the efficacy of drugsZhou et al. (2018). Chemical reactions in the solid state, as opposed to synthesis in solution, can open up novel reaction pathways and selectivitiesGalwey (2014). However, characterizing these transformations is challenging due to the complex interplay between molecular reorganization and the constraints imposed by the crystal lattice.
@@
The primary computational approach to investigating reactions or transformations involves identifying the minimum-energy pathway (or reaction coordinate) connecting the initial and final states of a transformation. This is typically achieved through the use of the nudged elastic band (NEB) methodJónsson, Mills, and Jacobsen (1998); Henkelman, Uberuaga, and Jónsson (2000); Sheppard et al. (2012); Mader et al. (2018); Shin et al. (2019); Han and Luber (2020); Hossain et al. (2020); Podewitz, Sen, and Buchmeiser (2021), which efficiently samples the minimum-energy path by connecting a series of images or configurations along the reaction path. NEB is particularly effective for determining transition states and energy barriers, as it maintains a series of replicas of the system that are connected by springs, ensuring smooth transitions along the reaction coordinate. Therefore, this constitutes the main methodology for studying chemical reactions in vacuum or within implicit solvent models.
@@
However, NEB calculations for solid-state systems are often hindered by the complexity of generating realistic initial reaction pathways, particularly for periodic molecular crystal structures. Linear interpolation, the simplest method for generating these paths, often fails to capture the intricate changes in both molecular orientation and unit cell parameters that occur during solid-state transformations, frequently leading to atomic collisions or nonphysical geometries. While in isolated systems problems with the interpolation can be corrected by hand, such manual corrections are extremely cumbersome for complex periodic structures.
@@
Another promising method for modeling dynamic processes is molecular dynamics (MD), which has been applied to polymorphic transitions of molecular crystals utilizing traditional force fieldsSalvalaglio et al. (2014); Gimondi and Salvalaglio (2017); Francia, Price, and Salvalaglio (2021). However, most traditional force fields are not designed to handle the breaking of chemical bonds, which needs to occur in chemical reactions, and their accuracy is often not sufficient. The main workhorse for the modeling of molecular crystalsOtero-de-la-Roza and Johnson (2012); Reilly and Tkatchenko (2013); Marom et al. (2013); Reilly and Tkatchenko (2015); Brandenburg and Grimme (2016); Shtukenberg et al. (2017); Hoja et al. (2019); Dolgonos, Hoja, and Boese (2019); Jana et al. (2021); Price, Otero-de-la-Roza, and Johnson (2023); Hoja, List, and Boese (2023) is density functional theory (DFT), since it provides a good balance between accuracy and computational cost. While first-principles MD simulations can be performed for small molecules in vacuum, modeling the kinetics of polymorph transitions and solid-state reactions with either MD or NEB for realistic systems is typically out of reach in terms of available computation times. In recent years, huge advancements have been made in the field of machine-learned force fields (MLFF), especially in the development of neural networksSchütt et al. (2017); Schütt, Unke, and Gastegger (2021); Unke et al. (2021); Batzner et al. (2022); Batatia et al. (2022); Frank, Unke, and Müller (2022); Frank et al. (2024). With an appropriate set of accurate reference data — typically high-level DFT calculations — MLFFs can be trained so that they are almost as efficient as traditional force fields but also much more accurate for specific applications.
@@
In this paper, we present a novel interpolation method to generate high-quality initial reaction paths for NEB calculations. The proposed method is designed to handle both rigid and flexible molecular systems, automatically avoiding atomic collisions and ensuring smooth transitions between the initial and final configuration. Furthermore, we also benchmark the applicability and accuracy of state-of-the-art MLFFs in the context of periodic NEB calculations. For that we generate tailor-made DFT reference data and utilized the newly developed equivariant message-passing neural network architecture SO3kratesFrank, Unke, and Müller (2022); Frank et al. (2024).
@@
This methodology is then tested on the polymorph transitions of the (low-temperature) polymorph of 1,4-diiodobenzeneAlcobé et al. (1994) into the (high-temperature) polymorph as well as that of theophylline polymorph I Khamar et al. (2011) into polymorph II.Fucke et al. (2012a) In addition, we have also modeled the solid-state reaction of a cocrystal composed of bis(N-allylimino)-1,4-dithiin and 9-bromoanthracene into the Diels-Alder productKhorasani et al. (2015). We show that our new interpolation method automatically produces viable initial reaction paths and that the obtained MLFFs are providing very accurate results compared to the much more expensive DFT reference data.
@@ II Computational Methods
For the three considered transitions, we have utilized the following experimentally determined crystal structures as starting point, with the Cambridge structural database code listed in parenthesis: 1,4-diiodobenzene polymorph (ZZZPRO05), 1,4-diiodobenzene polymorph (ZZZPRO07), theophylline polymorph I (BAPLOT04), theophylline polymorph II (BAPLOT06), co-crystal of bis(N-allylimino)-1,4-dithiin and 9-bromoanthracene (FUSCIH), and the respective Diels-Alder product (FURROB). All these structures are visualized in Fig. 1.
First, all structures were fully optimized with PBE+MBD-NLPerdew, Burke, and Ernzerhof (1996); Tkatchenko et al. (2012); Ambrosetti et al. (2014); Hermann and Tkatchenko (2020) utilizing the 2020 light species default settings for basis functions and integration grids within FHI-aimsBlum et al. (2009); Knuth et al. (2015); Ren et al. (2012); Yu et al. (2018); Havu et al. (2009); Ihrig et al. (2015) (version 231212). The k-point grids for the DFT part and the MBD part were chosen so that the number of k-points Å or 30 Å, respectively, where is the cell length in the respective direction. We have applied the following convergence criteria: eV for the total energy, 10-5 electrons per Å3 for the charge density, 10-4 eV/Å for forces, 10-3 eV for the sum of the eigenvalues, and 0.005 eV/Å during geometry optimizations. All calculations were carried out via the Atomic Simulation Environment ASELarsen et al. (2017).
Since a simple linear interpolation does not allow the generation of reasonable starting structures for a reaction path sampling via nudged elastic band (NEB) calculations for all investigated systems, we have developed the following interpolation methodology. Our approach to NEB pathway generation introduces a hybrid interpolation method which uses a linear interpolation for the unit cell and leverages the strengths of the spherical linear interpolation algorithmShoemake (1985) (SLERP) for molecular structures
with and being the quaternions representing the initial and final structures, and being the interpolation parameter between 0 and 1. SLERP was for instance already utilized for protein calculationsNguyen, Jaillet, and Redon (2017, 2018).
@@
Since these fragment interpolations can lead to atomic collision in the unit cell for some images, we employ a post-processing step that adjusts the unit cell size to ensure chemically reasonable structures throughout the pathway. Therefore, we first detect possible atomic collisions by searching across all images for intermolecular distances smaller than times the respective covalent radius as used by ASE. In the image with the maximum number of collisions we obtain the smallest occurring intermolecular distance. Then, the unit cell vectors for that particular image are increased by the the respective projection of the identified distance vector. Then the linear interpolation of the unit cell vectors is repeated for the whole pathway utilizing the newly changed cell parameters as fixed intermediate image. This process is repeated until the resulting pathway is completely free of collisions.
@@
This initial interpolation methodology was utilized throughout. For 1,4-diiodobenzene we have additionally also used a simple linear interpolation for comparison, for the other systems this leads to atomic collisions.
@@
Next, we have performed nudged elastic band (NEB) calculations utilizing the solid-state NEBSheppard et al. (2012) as implemented in TSASE, which allows for variable unit cells during the reaction path sampling. We always applied the climbing image methodHenkelman, Uberuaga, and Jónsson (2000) and used as convergence criteria 0.005 eV/Å for 1,4-diiodobenzene and 0.1 eV/Å for the two remaining, more complex systems.
@@
For 1,4-diiodobenzene we have performed NEB calculations with PBE+MBD utilizing the same settings as for the lattice relaxations described above. In this case, we have also performed a harmonic phonon calculation for the obtained transition-state estimate to validate the NEB results. For this calculation, we applied PBE+MBD utilizing phonopyTogo and Tanaka (2015) together with FHI-aims and a supercell of at least 12 Å was used to avoid artifacts. Furthermore, we have also performed a NEB pathway optimization for 1,4-diiodobenzene using DFTB3+MBD as implemented in DFTB+Hourahine et al. (2020), with a k-point grid satisfying Å.
@@
Next, we have created PBE+MBD reference data for the training of machine-learned force fields (MLFF) suitable for NEB calculations. Therefore, we have performed for all systems 5 ps NPT molecular dynamics (MD) simulations at 500 K utilizing FHI-aims together with ASE. We have used a timestep of 1 fs, a characteristic timescale of 15 fs for the thermostat, and 2000 fs2 for the barostat constant (pfactor). These simulations were performed at ambient pressure as well as at 1 GPa. For 1,4-diiodobenzene and theophylline we have utilized a 1x2x2 and 1x3x1 supercell, respectively. In addition, we have performed full lattice relaxations of all initial reaction-path images directly after interpolation. In this way, the reference energies and forces should contain a sufficiently large portion of the potential energy surface, suitable for the specific NEB calculations.
@@
The obtained PBE+MBD data was now used to train SO3kratesFrank, Unke, and Müller (2022); Frank et al. (2024) MLFFs for 2000 epochs, utilizing mostly default settings. For 1,4-diiodobenzene and theophylline we utilized 5000 samples for training and for the solid-state reaction 4000, while 1000 samples were always used for validation. Subsequently, lattice relaxations and subsequent NEB calculations were carried out for all systems in the same fashion as described for the PBE+MBD calculations above. We have utilized the GLPLanger, Frank, and Knoop (2023) package to obtain ASE calculators of our three MLFFs.
@@
Finally, we have performed several single-point calculations on top of the obtained optimized NEB reaction paths. On top of the SO3krates structures, we have performed PBE+MBD/light single points within FHI-aims as described above. For the PBE+MBD/light reaction path of 1,4-diiodbenzene, we have additionally also calculated PBE+MBD energies utilizing tight species default settings and ME3(PBE0+MBD:PBE+MBD)/tight multimer embeddingHoja, List, and Boese (2023) energies, essentially approximating tight PBE0+MBD results.

### v2

@@ Abstract
The modeling of solid-state transformations, such as polymorphic transitions and chemical reactions in molecular crystals, is vital for many applications including drug design or the development of new synthesis methods. However, a description via nudged-elastic band (NEB) calculations faces several crucial challenges. First, the automatic initial pathway generation based on a linear interpolation often fails for periodic systems, leading to unrealistic geometries and atomic collisions. Second, the necessary system sizes are typically beyond the scope of density functional theory (DFT) calculations in terms of computational cost, but the associated accuracy is vitally needed. To address these issues, we introduce a hybrid interpolation method that combines linear interpolation for cell parameters with spherical linear interpolation (SLERP) for molecular structures or intramolecular fragments, ensuring smooth and realistic transitions. Moreover, we train and benchmark machine-learned force fields (MLFFs) based on the SO3krates equivariant neural network architecture to accelerate NEB calculations while retaining near-DFT accuracy. We apply our approach to two polymorph transitions and a solid-state Diels-Alder reaction and show that our new interpolation method reliably produces viable initial pathways. The MLFFs are trained on PBE+MBD reference data and reproduce DFT lattice energies with a mean absolute error of 0.4 kJ mol-1 along the minimum-energy paths. These results highlight the potential of combining advanced interpolation techniques with MLFFs to enable automated, accurate, and efficient exploration of solid-state transformations in molecular crystals.
@@ I Introduction
Understanding solid-state transformations, such as transitions between molecular crystal polymorphs or solid-state chemical reactions at a molecular level is vital for pharmaceuticals, chemistry, and materials scienceKaupp (2002); Datta and Grant (2004); Anthony (2006); Saha, Nath, and Thakuria (2022). Polymorphs are different crystal-packing arrangements of the same molecule, and although the chemical composition of a molecular crystal does not change when one polymorph transforms into another, this can often be associated with quite drastic changes in several properties.
An important issue for the pharmaceutical industry is that polymorph transitions can affect the efficacy of drugsZhou et al. (2018), exemplified by the HIV medication ritonavirBauer et al. (2001); Bučar, Lancaster, and Bernstein (2015) that became unusable after an unexpected formation of a less soluble polymorph. Therefore, crystal structure predictionPrice (2014) (CSP), which yields a polymorphic energy landscape of a given molecular crystal, has become an invaluable tool for gaining knowledge about polymorphism and the accuracy of computational approaches is regularly assessed by blind tests organized by the Cambridge Crystallographic Data CentreReilly et al. (2016); Hunnisett et al. (2024a, b). The resulting thermodynamical polymorph stability rankings often include many more structures than are found experimentally since kinetic factors are not taken into accountPrice (2013, 2022); Beran (2023). Therefore, understanding the kinetics of polymorph transitions is key to solving this overprediction problem. Moreover, studying chemical reactions in the solid state is becoming ever more important since these reactions, as opposed to synthesis in solution, can open up novel reaction pathways and selectivitiesGalwey (2014). However, characterizing these transformations is challenging due to the complex interplay between molecular reorganization and the constraints imposed by the crystal lattice.
One key method for modeling solid-state transformations in molecular crystals is molecular dynamics (MD) but the relevant transitions are rarely observed in unbiased trajectories, so enhanced-sampling schemes are usually indispensable. Several metadynamics studiesSalvalaglio et al. (2014); Gimondi and Salvalaglio (2017); Francia et al. (2020); Francia, Price, and Salvalaglio (2021) focused on polymorph transitions of molecular crystals utilizing traditional force fields and could reduce the number of structures in CSP landscapes to a manageable set of distinct finite-temperature basins that better reflect experimental reality. Another promising methodology for reducing the overprediction of structures in molecular CSP is the threshold algorithmYang and Day (2022); Butler and Day (2023) based on Monte Carlo simulations, which provides a mapping of connected structures into basins and also yields approximate inter-polymorph barriers. However, most traditional force fields are not designed to handle the breaking of chemical bonds, which needs to occur when modeling chemical reactions, and their accuracy is often not sufficient when we are interested in the actual energy barriers (or reaction rates) of a transformation.
@@
The primary computational approach to investigate the barriers of reactions or transformations on a first-principles level involves identifying the minimum-energy pathway (or reaction coordinate) that connects the initial and final states of a transformation. This is typically achieved through the use of the nudged elastic band (NEB) method or similar approachesJónsson, Mills, and Jacobsen (1998); Henkelman, Uberuaga, and Jónsson (2000); Sheppard et al. (2012); Mader et al. (2018); Shin et al. (2019); Han and Luber (2020); Hossain et al. (2020); Podewitz, Sen, and Buchmeiser (2021), which efficiently samples the minimum-energy path by connecting a series of images or configurations along a reaction path. NEB is particularly effective for determining transition states and energy barriers, as it maintains a series of replicas of the system that are connected by springs, ensuring smooth transitions along the reaction coordinate. Therefore, this constitutes the main methodology for studying chemical reactions in vacuum or within implicit solvent models.
@@
However, NEB calculations for solid-state systems are often hindered by the complexity of generating realistic initial reaction pathways, particularly for periodic molecular crystal structures. Linear interpolation, the simplest method for generating these paths, often fails to capture the intricate changes in both molecular orientation and unit cell parameters that occur during solid-state transformations, frequently leading to atomic collisions or nonphysical geometries. Although in isolated systems problems with the interpolation can be corrected by hand, such manual corrections are extremely cumbersome for complex periodic structures.
@@
The main workhorse for the modeling of molecular crystals on a first-principles levelOtero-de-la-Roza and Johnson (2012); Reilly and Tkatchenko (2013); Marom et al. (2013); Reilly and Tkatchenko (2015); Brandenburg and Grimme (2016); Shtukenberg et al. (2017); Hoja et al. (2019); Dolgonos, Hoja, and Boese (2019); Jana et al. (2021); Price, Otero-de-la-Roza, and Johnson (2023); Hoja, List, and Boese (2023) is density functional theory (DFT), since it provides a good balance between accuracy and computational cost. While first-principles MD simulations can be performed for small molecules in vacuum, modeling the kinetics of polymorph transitions and solid-state reactions with either MD or NEB for realistic systems is typically out of reach in terms of available computation times. In recent years, great advances have been made in the field of machine-learned force fieldsChmiela et al. (2017); Schütt et al. (2017); Chmiela et al. (2018); Schütt, Unke, and Gastegger (2021); Poltavsky and Tkatchenko (2021); Unke et al. (2021); Sauceda et al. (2022); Batzner et al. (2022); Batatia et al. (2022); Frank, Unke, and Müller (2022); Frank et al. (2024); Kovács et al. (2025) (MLFFs), especially in the development of neural networks, and these methods are increasingly utilized for molecular crystalsRaimbault et al. (2019); Krynski and Rossi (2021); Lewis, Lazzaroni, and Rossi (2023); Žugec, Geilhufe, and Lončarić (2024); Mladineo and Lončarić (2024); Zhao et al. (2025); Nickerson and Johnson (2025). With an appropriate set of accurate reference data — typically high-level DFT calculations — MLFFs can be trained so that they are almost as efficient as traditional force fields but also much more accurate for specific applications.
@@
In this paper, we illustrate how energy barriers of molecular crystal transformations can be obtained with near-DFT accuracy relying solely on NEB calculations. First, we present a novel interpolation method to generate high-quality initial reaction paths for NEB calculations of molecular crystals. The proposed method is designed to handle both rigid and flexible molecular systems, automatically avoiding atomic collisions and ensuring smooth transitions between the initial and final configurations. Then, we also train and benchmark the applicability and accuracy of state-of-the-art MLFFs in the context of periodic NEB calculations. For that we generate tailor-made DFT reference data and utilize the equivariant message-passing neural network architecture SO3kratesFrank, Unke, and Müller (2022); Frank et al. (2024).
@@
This methodology is then tested on the polymorph transitions of the (low-temperature) polymorph of 1,4-diiodobenzeneAlcobé et al. (1994) into the (high-temperature) polymorph as well as that of theophylline polymorph I Khamar et al. (2011) into polymorph II.Fucke et al. (2012a) In addition, we have also modeled the solid-state reaction of a cocrystal composed of bis(N-allylimino)-1,4-dithiin and 9-bromoanthracene into the Diels-Alder productKhorasani et al. (2015). We show that our new interpolation method automatically produces viable initial reaction paths and that the obtained MLFFs are providing very accurate results compared to the much more expensive DFT reference data for energy barriers of polymorph transitions as well as solid-state reactions.
@@ II Computational Methods
For the three considered transitions, we have utilized the following experimentally determined crystal structures as starting points, with the Cambridge structural database code listed in parenthesis: 1,4-diiodobenzene polymorph (ZZZPRO05), 1,4-diiodobenzene polymorph (ZZZPRO07), theophylline polymorph I (BAPLOT04), theophylline polymorph II (BAPLOT06), co-crystal of bis(N-allylimino)-1,4-dithiin and 9-bromoanthracene (FUSCIH), and the respective Diels-Alder product (FURROB). All these structures are visualized in Fig. 1.
First, all structures were fully optimized with PBE+MBD-NLPerdew, Burke, and Ernzerhof (1996); Tkatchenko et al. (2012); Ambrosetti et al. (2014); Hermann and Tkatchenko (2020) utilizing the 2020 light species default settings for basis functions and integration grids within FHI-aimsBlum et al. (2009); Knuth et al. (2015); Ren et al. (2012); Yu et al. (2018); Havu et al. (2009); Ihrig et al. (2015) (version 231212). The k-point grids for the DFT part and the MBD part were chosen so that the number of k-points Å or 30 Å, respectively, where is the cell length in the respective direction. We have applied the following convergence criteria: eV for the total energy, 10-5 electrons per Å3 for the charge density, 10-4 eV Å-1 for forces, 10-3 eV for the sum of the eigenvalues, and 0.005 eV Å-1 during geometry optimizations for the maximal force component. All calculations were carried out via the Atomic Simulation Environment ASELarsen et al. (2017).
Since a simple linear interpolation does not allow the generation of reasonable starting structures for a reaction path sampling via nudged elastic band (NEB) calculations for all investigated systems, we have developed the following interpolation methodology. Our approach to NEB pathway generation introduces a hybrid interpolation method, which uses a linear interpolation for the unit cell and leverages the strengths of the spherical linear interpolation algorithmShoemake (1985) (SLERP) for molecular structures
with and being the quaternions representing the initial and final structures, and being the interpolation parameter between 0 and 1. SLERP was, for example, already utilized for protein calculationsNguyen, Jaillet, and Redon (2017, 2018).
@@
Since these fragment interpolations can lead to atomic collision in the unit cell for some images, we employ a post-processing step that adjusts the unit cell size to ensure chemically reasonable structures throughout the pathway. Therefore, we first detect possible atomic collisions by searching across all images for intermolecular distances smaller than times the respective covalent radius as used by ASE. In the image with the maximum number of collisions we obtain the smallest occurring intermolecular distance. Then, the unit cell vectors for that particular image are increased by the respective projection of the identified distance vector. Next, the linear interpolation of the unit cell vectors is repeated for the whole pathway utilizing the newly changed cell parameters as fixed intermediate image. This process is repeated until the resulting pathway is completely free of collisions.
@@
This initial interpolation methodology was utilized throughout. For 1,4-diiodobenzene we have additionally used a simple linear interpolation for comparison, for the other systems this leads to atomic collisions.
@@
Next, we have performed nudged elastic band (NEB) calculations utilizing the solid-state NEB methodSheppard et al. (2012) as implemented in TSASE, which allows for variable unit cells during the reaction path sampling. We always applied the climbing image methodHenkelman, Uberuaga, and Jónsson (2000) and used as convergence criteria 0.005 eV Å-1 for 1,4-diiodobenzene and 0.1 eV Å-1 for the two remaining, more complex systems.
@@
For 1,4-diiodobenzene we have performed NEB calculations with PBE+MBD utilizing the same settings as for the lattice relaxations described above. In this case, we have also performed a harmonic phonon calculation for the obtained transition-state estimate to validate the NEB results, as well as for the optimized structures of both polymorphs. For this calculation, we applied PBE+MBD utilizing phonopyTogo and Tanaka (2015) together with FHI-aims using a supercell of at least 12 Å in each direction to avoid artifacts and finite displacements of 0.005 Å. The q-grid for the evaluation of vibrational free energies was chosen so that the number of q-points () in every direction satisfies Å, with being the unit-cell length in the respective direction. Furthermore, we have also performed a NEB pathway optimization for 1,4-diiodobenzene using DFTB3+MBD as implemented in DFTB+Hourahine et al. (2020), with a k-point grid satisfying Å and utilizing the 3ob parameter setGaus, Goez, and Elstner (2012); Gaus et al. (2014); Kubillus et al. (2014).
@@
Next, we have created PBE+MBD reference data for the training of machine-learned force fields (MLFFs) suitable for the NEB calculations of interest. Therefore, we have performed for all systems several short 5 ps molecular dynamics (MD) simulations with the isothermal-isobaric ensemble (NPT) at 500 K utilizing FHI-aims together with ASE. We have used a timestep of 1 fs and characteristic timescales of 15 fs and 45 fs for the thermostat and the barostat constant, respectively. For every system, we have performed separate 5 ps simulations for the initial and final structures as well as for a structure directly in between according to our introduced interpolation. Furthermore, all these simulations were not only performed at ambient pressure but also at a pressure of 1 GPa. We note that the goal of these MD simulations is not to obtain proper dynamics of the systems but rather sample a diverse set of structures close to the ones observed along the respective transition paths we aim to study. For 1,4-diiodobenzene and theophylline we have utilized a 1x2x2 and 1x3x1 supercell, respectively.
In addition, we have performed full lattice relaxations of all initial reaction-path images directly after interpolation. In this way, the collective reference energies and forces from several MD simulations and optimization trajectories should cover a sufficiently large portion of the needed potential energy surface, suitable for the specific NEB calculations. In order to remove duplicate structures and high-energy structures we have ensured that the final reference data only include structures that have a RMSD of at least 0.04 Å w.r.t. all other structures and from the individual MD trajectories we have removed all structures, which were by more than 1.5 standard deviations less stable than the average energy of the respective trajectory.
@@
The obtained PBE+MBD data was now used to train for each system an individual SO3kratesFrank, Unke, and Müller (2022); Frank et al. (2024) MLFF model with a local neighborhood cutoff of 5 Å, applied minimal image convention, a feature dimension of 132, 3 layers, 4 heads in the message-passing update, and degrees in the Euclidean variables for 2000 epochs. For all three models, we utilized 5000 samples for training and 1000 samples for validation. Subsequently, lattice relaxations and subsequent NEB calculations were carried out for all systems in the same fashion as described for the PBE+MBD calculations above. We have utilized the GLPLanger, Frank, and Knoop (2023) package to obtain ASE calculators of our three MLFFs.
@@
Finally, we have performed several single-point calculations on top of the obtained optimized NEB reaction paths, starting with PBE+MBD/light single points within FHI-aims (as described above) using the SO3krates structures. For the PBE+MBD/light reaction path of 1,4-diiodobenzene, we have additionally calculated PBE+MBD energies utilizing tight species default settings and ME3(PBE0+MBD:PBE+MBD)/tight multimer embeddingHoja, List, and Boese (2023) energies, essentially approximating tight PBE0+MBD results.